# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ D.1 Main Proof of Theorem 1
$\leq\frac{2\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left(\frac{1}{|S_{t}|}\right)^{2}|S_{t}|}\sqrt{\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi(x_{t},\bar{a}_{t})\right\|_{H_{t}^{-1}}^{2}}$ (Cauchy-Schwartz ineq.)
@@
$=\frac{2\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\left(\frac{1}{|S_{t}|}\right)^{2}|S_{t}|}\sqrt{50\sum_{t\notin{\mathcal{T}}_{0}\cup{\mathcal{T}}^{w}}\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi(x_{t},\bar{a}_{t})\right\|_{\Lambda_{t}^{-1}}^{2}}$ (Lemma D.2, with prob. $1-\delta$)
@@
$\leq\frac{15\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t=1}^{T}\min\left\{1,\sum_{a\in S_{t}}\left\|\phi\left(x_{t},a\right)-\phi(x_{t},\bar{a}_{t})\right\|_{\Lambda_{t}^{-1}}^{2}\right\}}$
@@
$\leq\frac{15\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{2d\log\left(1+\frac{2KT}{d\lambda}\right)}$ (Lemma D.3)
By combining Equations (D.4), (D.5), and (D.7), and setting $\beta_{T+1}(\delta)=\mathcal{O}\big(B\sqrt{d\log(KT)}+B\sqrt{\lambda}\big)$, we derive that, with probability at least $1-3\delta$ (omitting logarithmic terms and polynomial dependencies on $B$ for brevity),
@@
$\textup{{SubOpt}}(T)$ $=\tilde{\mathcal{O}}\left(\frac{d}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}+\frac{d^{2}K^{2}}{\kappa T}\right).$
@@
D.2 Proofs of Lemmas for Theorem 1
D.2.1 Proof of Lemma D.2
Proof of Lemma D.2.
@@ D.2 Proofs of Lemmas for Theorem 1
where each row corresponds to the feature vector of an action ranked from position $j$ to $|S_{s}|$ in the ranking $\sigma_{s}$. Moreover, we define $a^{\star}_{sj}=\phi(x_{s},\sigma_{sj})^{\top}\bm{\theta}^{\star}$, $\mathbf{a}^{\star,(j)}_{s}=(a^{\star}_{sk})_{k=j}^{|S_{s}|}\in\mathbb{R}^{|S_{s}|-j+1}$.
Then, using the $\ell_{\infty}$-norm self-concordant property of the MNL loss [43], for any $s\in[t-1]\setminus{\mathcal{T}}^{w}$, we obtain
$\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}^{(j+1)}_{s})$ $=\left(\bm{\Phi}_{s}^{(j)}\right)^{\top}\nabla_{\mathbf{a}}^{2}\,\bar{\ell}^{(j)}_{s}(\mathbf{a}_{s}^{(j)})\,\bm{\Phi}_{s}^{(j)}$ (Eqn. (D.8))
$\succeq e^{-3\sqrt{2}\|\mathbf{a}_{s}^{(j)}-\mathbf{a}_{s}^{\star,(j)}\|_{\infty}}\left(\bm{\Phi}_{s}^{(j)}\right)^{\top}\nabla_{\mathbf{a}}^{2}\,\bar{\ell}^{(j)}_{s}(\mathbf{a}_{s}^{\star,(j)})\,\bm{\Phi}_{s}^{(j)}$ (Lemma F.1)
$\succeq\frac{1}{e}\left(\bm{\Phi}_{s}^{(j)}\right)^{\top}\nabla_{\mathbf{a}}^{2}\,\bar{\ell}^{(j)}_{s}(\mathbf{a}_{s}^{\star,(j)})\,\bm{\Phi}_{s}^{(j)}$ ($\|\mathbf{a}_{s}^{(j)}-\mathbf{a}_{s}^{\star,(j)}\|_{\infty}\leq\frac{1}{3\sqrt{2}}$)
$=\frac{1}{e}\nabla^{2}\ell^{(j)}_{s}(\bm{\theta}^{\star}),$ (Eqn. (D.8))
@@
$\|\mathbf{a}_{s}^{(j)}-\mathbf{a}_{s}^{\star,(j)}\|_{\infty}$ $=\max_{k=j,\dots,|S_{s}|}\left|\phi(x_{k},\sigma_{sk})^{\top}\left(\widehat{\bm{\theta}}_{s}^{(k+1)}-\bm{\theta}^{\star}\right)\right|$
$\leq\max_{k=j,\dots,|S_{s}|}\left\|\phi(x_{k},\sigma_{sk})\right\|_{H_{s}^{-1}}\left\|\widehat{\bm{\theta}}_{s}^{(k+1)}-\bm{\theta}^{\star}\right\|_{H_{s}}$ (Hölder’s inequality)
$\leq\frac{1}{3\sqrt{2}\beta_{T+1}(\delta)}\max_{k=j,\dots,|S_{s}|}\left\|\widehat{\bm{\theta}}_{s}^{(k+1)}-\bm{\theta}^{\star}\right\|_{H_{s}^{(k+1)}}$ ($s\notin{\mathcal{T}}^{w}$, $H_{s}\preceq H_{s}^{(k+1)}$)
$\leq\frac{\beta_{T+1}(\delta)}{3\sqrt{2}\beta_{T+1}(\delta)}$ (Corollary D.1, $\beta_{t}(\delta)$ is non-decreasing)
$=\frac{1}{3\sqrt{2}}.$
@@
$H_{t}\succeq\!\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}^{(j+1)}_{s})+\lambda\mathbf{I}_{d}\succeq\frac{1}{e}\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s}(\bm{\theta}^{\star})+\lambda\mathbf{I}_{d}.$ (D.9)
Now, for better presentation, we define the Multinomial Logit (MNL) choice probability [48] for a given assortment $S$ at round $s$ as follows:
@@
Let $S_{s}=\{\sigma_{s1},\dots,\sigma_{s|S_{s}|}\}$. Thus, the PL model in Equation (1) can be rewritten as follows:
$\mathbb{P}(\sigma_{s}|x_{s},S_{s};\bm{\theta})$ $=P_{s}(\sigma_{s1}|S_{s};\bm{\theta})\cdot P_{s}(\sigma_{s2}|S_{s}\setminus\{\sigma_{s1}\};\bm{\theta})\cdot\;\dots\;\cdot P_{s}(\sigma_{s|S_{s}|}|\{\sigma_{s|S_{s}|}\};\bm{\theta})$
$=\prod_{j=1}^{|S_{s}|}P_{s}\left(\sigma_{sj}|\{\sigma_{sj},\dots,\sigma_{s|S_{s}|}\};\bm{\theta}\right).$
For simplicity, we define $S^{(j)}_{s}:=\{\sigma_{sj},\dots,\sigma_{s|S_{s}|}\}$, and let $P_{s}^{(j)}$ denote the (true) MNL distribution over the remaining actions in $S_{s}$ after removing the first j1j-1 selected actions, i.e., $P_{s}^{(j)}=P_{s}(\cdot|S^{(j)}_{s};\bm{\theta}^{\star})$. Then, to further lower bound the right-hand side of Equation (D.9), we proceed as follows:

### v2

@@ C.1 Main Proof of Theorem 1
$\quad\quad\leq\left(1+e^{\frac{\beta_{T+1}(\delta)}{3\sqrt{2}K\beta_{T+1}(\delta)}}\right)^{2}$ ($t\notin{\mathcal{T}}^{w}$ and Corollary C.1)
$\quad\quad=(1+e^{1/K})^{2}\leq(1+e)^{2}.$
Recall that $S^{(j)}_{\sigma}:=\{\sigma_{j},\dots,\sigma_{|S|}\}$ denotes the subset of remaining actions in $S$ after removing the first j1j-1 actions, given $S$ and $\sigma$. We denote the Multinomial Logit (MNL) model at round $t$ by $P_{t}(a\mid S;\bm{\theta})$ (see Equation (9)). Then, by our assortment selection rule in Equation (11), we obtain
$\dot{\mu}$ $\left(\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\bm{\theta}^{\star}\right)\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
$\leq\dot{\mu}\left(\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\widehat{\bm{\theta}}_{t}\right)e^{\left|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)^{\top}\big(\widehat{\bm{\theta}}_{t}-\bm{\theta}^{\star}\big)\right|}$
$\quad\quad\cdot\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$ (Lemma E.2)
$\leq e\cdot\dot{\mu}\left(\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$ (Eqn. C.5)
$=\frac{e}{2}\cdot\mathbb{E}_{a\sim P_{t}(\cdot|S^{\star}_{t};\widehat{\bm{\theta}}_{t})}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{a^{\prime}\sim P_{t}(\cdot|S^{\star}_{t};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{H_{t}^{-1}}\right]$ ($\dot{\mu}(z)=\mu(z)(1-\mu(z))$ and let $S^{\star}_{t}:=\{\pi^{\star}(x_{t}),\widehat{\pi}_{T}(x_{t})\}$)
$=\frac{e}{|S^{\star}_{t}|}\mathbb{E}_{\sigma\sim\mathbb{P}_{t}(\cdot|S^{\star}_{t};\widehat{\bm{\theta}}_{t})}\sum_{j=1}^{|S^{\star}_{t}|}\mathbb{E}_{a\sim P_{t}(\cdot|S^{\star,(j)}_{t};\widehat{\bm{\theta}}_{t})}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{a^{\prime}\sim P_{t}(\cdot|S^{\star,(j)}_{t};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{H_{t}^{-1}}\right],$ ($|S^{\star}_{t}|=2$)
where the second inequality holds since, for any $t\notin{\mathcal{T}}^{w}$ and $a,a^{\prime}\in\mathcal{A}$, the following property holds:
$\Big|\big($ $\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\left(\widehat{\bm{\theta}}_{t}-\bm{\theta}^{\star}\right)\Big|$
$\leq\big\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big\|_{H_{t}^{-1}}\big\|\widehat{\bm{\theta}}_{t}-\bm{\theta}^{\star}\big\|_{H_{t}}$ (Hölder’s inequality)
$\leq\frac{1}{3\sqrt{2}K\beta_{T+1}(\delta)}\big\|\widehat{\bm{\theta}}_{t}-\bm{\theta}^{\star}\big\|_{H_{t}}$ ($t\neq{\mathcal{T}}^{w}$)
$\leq\frac{\beta_{t}(\delta)}{K\beta_{T+1}(\delta)}$ (Corollary C.1)
$\leq\frac{1}{K}.$ (C.5)
Combining the above results, we get
$\sum_{t\notin{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
$\leq e(1+e)^{2}\sum_{t\notin{\mathcal{T}}^{w}}\frac{1}{|S^{\star}_{t}|}\sum_{j=1}^{|S^{\star}_{t}|}\mathbb{E}_{\begin{subarray}{c}\mathbb{P}_{t}(\cdot|S^{\star}_{t};\widehat{\bm{\theta}}_{t})\\ P_{t}(\cdot|S^{\star,(j)}_{t};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{P_{t}(\cdot|S^{\star,(j)}_{t};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{H_{t}^{-1}}\right]$ (C.6)
$\leq e(1+e)^{2}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\frac{|S_{t}|}{|S^{\star}_{t}|}\sum_{j=1}^{|S^{\star}_{t}|}\mathbb{E}_{\begin{subarray}{c}\mathbb{P}_{t}(\cdot|S^{\star}_{t};\widehat{\bm{\theta}}_{t})\\ P_{t}(\cdot|S^{\star,(j)}_{t};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{P_{t}(\cdot|S^{\star,(j)}_{t};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{H_{t}^{-1}}^{2}\right]}$ (Cauchy-Schwartz ineq., and ${\mathcal{T}}^{w}\subseteq[T]$)
$\leq e(1+e)^{2}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\frac{\cancel{|S_{t}|}}{\cancel{|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S_{t}}|}}\sum_{j=1}^{|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S_{t}}|}\mathbb{E}_{\begin{subarray}{c}\mathbb{P}_{t}(\cdot|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S_{t}};\widehat{\bm{\theta}}_{t})\\ P_{t}(\cdot|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S^{(j)}_{t}};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{P_{t}(\cdot|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S^{(j)}_{t}};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{H_{t}^{-1}}^{2}\right]}$ (Assortment selection rule, Equation (11))
$\leq e^{2}(1+e)^{2}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\begin{subarray}{c}\mathbb{P}_{t}(\cdot|S_{t};{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{\theta}^{\star}})\\ P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{H_{t}^{-1}}^{2}\right]}$ (Eqn. (C.7))
@@
$\leq\sqrt{3}e^{3}(1+e)^{2}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\begin{subarray}{c}\mathbb{P}_{t}(\cdot|S_{t};\bm{\theta}^{\star})\\ P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}M_{t}^{-1}}}^{2}\right]},$ (Lemma C.2, with prob. $1-\delta$)
where the second-to-last inequality holds because, for any $t\notin{\mathcal{T}}^{w}$, $S\in{\mathcal{S}}$, and $a\in S$, we have
$P_{t}(a|S;\widehat{\bm{\theta}}_{t})$ $=\frac{\exp\left(\phi(x_{t},a)^{\top}\widehat{\bm{\theta}}_{t}\right)}{\sum_{a^{\prime}\in S}\exp\left(\phi(x_{t},a^{\prime})^{\top}\widehat{\bm{\theta}}_{t}\right)}$
$=\frac{\exp\left(\big(\phi(x_{t},a)-\phi(x_{t},\bar{a})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)}{1+\sum_{a^{\prime}\in S\setminus\{\bar{a}\}}\exp\left(\big(\phi(x_{t},a^{\prime})-\phi(x_{t},\bar{a})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)}$ (any $\bar{a}\in S$)
$\leq\frac{\exp\left(\big(\phi(x_{t},a)-\phi(x_{t},\bar{a})\big)^{\top}\bm{\theta}^{\star}+1/K\right)}{1+\sum_{a^{\prime}\in S\setminus\{\bar{a}\}}\exp\left(\big(\phi(x_{t},a^{\prime})-\phi(x_{t},\bar{a})\big)^{\top}\bm{\theta}^{\star}-1/K\right)}$ (Eqn. (C.5))
$=e^{2/K}P_{t}(a|S;\bm{\theta}^{\star}),$
which implies that, for the ranking $\sigma=(\sigma_{1},\dots,\sigma_{|S|})$,
$\mathbb{P}_{t}(\sigma|S;\widehat{\bm{\theta}}_{t})$ $=\prod_{j=1}^{|S|}P_{t}(\sigma_{j}|S^{(j)}_{\sigma};\widehat{\bm{\theta}}_{t})\leq e^{2|S|/K}\prod_{j=1}^{|S|}P_{t}(\sigma_{j}|S^{(j)}_{\sigma};\bm{\theta}^{\star})\leq e^{2}\prod_{j=1}^{|S|}P_{t}(\sigma_{j}|S^{(j)}_{\sigma};\bm{\theta}^{\star})$
$=e^{2}\mathbb{P}_{t}(\sigma|S;\bm{\theta}^{\star}).$ (C.7)
Moreover, by setting $\lambda=d\log T$ and $T\geq e^{K/d}$, for any $t$, we have
$\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\begin{subarray}{c}\mathbb{P}_{t}(\cdot|S_{t};\bm{\theta}^{\star})\\ P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{M_{t}^{-1}}^{2}\right]$ $\leq\frac{2K}{\lambda}\leq 2.$ (C.8)
Hence, by the elliptical potential for expected mean-centered uncertainty (Lemma C.3), we obtain
$\frac{1}{T}\sum_{t\notin{\mathcal{T}}^{w}}\big(\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\big)^{\top}\left(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{T+1}\right)$
@@
$\leq\frac{\beta_{T+1}(\delta)}{T}\sum_{t\notin{\mathcal{T}}^{w}}\left\|\phi\left(x_{t},\pi^{\star}(x_{t})\right)-\phi\left(x_{t},\widehat{\pi}_{T}(x_{t})\right)\right\|_{H_{t}^{-1}}$
$\leq\mathcal{O}\!\left(\frac{\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{\sum_{t\notin{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\begin{subarray}{c}\mathbb{P}_{t}(\cdot|S_{t};\bm{\theta}^{\star})\\ P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{P_{t}(\cdot|S_{t}^{(j)};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{M_{t}^{-1}}^{2}\right]}\right)$
@@
$=\mathcal{O}\!\left(\frac{\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\sqrt{d\log\left(1+\frac{KT}{d\lambda}\right)}\right)$ (Lemma C.3 with Eqn. (C.8))
@@
$=\mathcal{O}\!\left(\frac{\beta_{T+1}(\delta)}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}\cdot\sqrt{d\log\left(KT\right)}\right).$ (C.9)
By combining Equations (C.3), (C.4), and (C.9), and setting $\beta_{T+1}(\delta)=\mathcal{O}\big(B\sqrt{d\log(KT)}+B\sqrt{\lambda}\big)$, we derive that, with probability at least $1-3\delta$ (omitting logarithmic terms and polynomial dependencies on $B$ for brevity),
$\textup{{SubOpt}}(T)$ $=\tilde{\mathcal{O}}\left(\frac{d}{T}\sqrt{\sum_{t=1}^{T}\frac{1}{|S_{t}|}}+\frac{d^{2}K^{4}}{\kappa T}\right).$
@@
C.2 Proofs of Lemmas for Theorem 1
C.2.1 Proof of Lemma C.2
Proof of Lemma C.2.
@@ C.2 Proofs of Lemmas for Theorem 1
where each row corresponds to the feature vector of an action ranked from position $j$ to $|S_{s}|$ in the ranking $\sigma_{s}$. Moreover, we define $\widehat{a}_{sj}=\phi(x_{s},\sigma_{sj})^{\top}\widehat{\bm{\theta}}_{s}$ and $\widehat{\mathbf{a}}^{(j)}_{s}=(\widehat{a}_{sk})_{k=j}^{|S_{s}|}\in\mathbb{R}^{|S_{s}|-j+1}$
Then, using the $\ell_{\infty}$-norm self-concordant property of the MNL loss [lee2025improved], for any $s\in[t-1]\setminus{\mathcal{T}}^{w}$, we obtain
$\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}^{(j+1)}_{s})$ $=\left(\bm{\Phi}_{s}^{(j)}\right)^{\top}\nabla_{\mathbf{a}}^{2}\,\bar{\ell}^{(j)}_{s}(\mathbf{a}_{s}^{(j)})\,\bm{\Phi}_{s}^{(j)}$ (Eqn. (C.10))
$\succeq e^{-3\sqrt{2}\|\mathbf{a}_{s}^{(j)}-\widehat{\mathbf{a}}_{s}^{(j)}\|_{\infty}}\left(\bm{\Phi}_{s}^{(j)}\right)^{\top}\nabla_{\mathbf{a}}^{2}\,\bar{\ell}^{(j)}_{s}(\widehat{\mathbf{a}}_{s}^{(j)})\,\bm{\Phi}_{s}^{(j)}$ (Lemma E.1)
$\succeq\frac{1}{e^{2}}\left(\bm{\Phi}_{s}^{(j)}\right)^{\top}\nabla_{\mathbf{a}}^{2}\,\bar{\ell}^{(j)}_{s}(\widehat{\mathbf{a}}_{s}^{(j)})\,\bm{\Phi}_{s}^{(j)}$ ($\|\mathbf{a}_{s}^{(j)}-\widehat{\mathbf{a}}_{s}^{(j)}\|_{\infty}\leq\frac{2}{3\sqrt{2}}$)
$=\frac{1}{e^{2}}\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}_{s}),$ (Eqn. (C.10))
@@
$\|\mathbf{a}_{s}^{(j)}-\widehat{\mathbf{a}}_{s}^{(j)}\|_{\infty}$ $\leq\max_{k=j,\dots,|S_{s}|}\left|\phi(x_{k},\sigma_{sk})^{\top}\left(\widehat{\bm{\theta}}_{s}^{(k+1)}-\widehat{\bm{\theta}}_{s}\right)\right|$
$\leq\max_{k=j,\dots,|S_{s}|}\left\|\phi(x_{k},\sigma_{sk})\right\|_{H_{s}^{-1}}\left(\left\|\widehat{\bm{\theta}}_{s}^{(k+1)}-\bm{\theta}^{\star}\right\|_{H_{s}}+\left\|\widehat{\bm{\theta}}_{s}-\bm{\theta}^{\star}\right\|_{H_{s}}\right)$ (Hölder’s inequality)
$\leq\frac{1}{3\sqrt{2}\beta_{T+1}(\delta)}\max_{k=j,\dots,|S_{s}|}\left(\left\|\widehat{\bm{\theta}}_{s}^{(k+1)}-\bm{\theta}^{\star}\right\|_{H_{s}^{(k+1)}}+\left\|\widehat{\bm{\theta}}_{s}-\bm{\theta}^{\star}\right\|_{H_{s}}\right)$ ($s\notin{\mathcal{T}}^{w}$, $H_{s}\preceq H_{s}^{(k+1)}$)
$\leq\frac{2\beta_{T+1}(\delta)}{3\sqrt{2}\beta_{T+1}(\delta)}$ (Corollary C.1, $\beta_{t}(\delta)$ is non-decreasing)
$=\frac{2}{3\sqrt{2}}.$
@@
$H_{t}\succeq\!\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}^{(j+1)}_{s})+\lambda\mathbf{I}_{d}\succeq\frac{1}{e^{2}}\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}_{s})+\lambda\mathbf{I}_{d}.$
Recall the definition of the MNL choice probability (9):
@@
Given $S_{s}=\{\sigma_{s1},\dots,\sigma_{s|S_{s}|}\}$ and the ranking feedback $\sigma_{s}$, we define $S^{(j)}_{\sigma_{s}}:=\{\sigma_{sj},\dots,\sigma_{s|S_{s}|}\}$ as the subset of remaining actions in $S_{s}$ after removing the first j1j-1 actions. For simplicity, let $\widehat{P}_{s}^{(j)}=P_{s}(\cdot|S^{(j)}_{\sigma_{s}};\widehat{\bm{\theta}}_{t})$. Then, we can express the Hessian of the loss as follows: